# AdS3 Integrability, Tensionless Limits, and Deformations: A Review

### v3

@@
Abstract
Motivated by the recent advances in the understanding of integrability for $AdS_{3}$ backgrounds, we present a lightning review of this approach, with particular attention to the “tensionless” limits (with zero and one unit of NSNS flux), and to the many integrable deformations of the supergravity backgrounds. Our aim is to concisely but comprehensively take stock of the state of the art in the field, in a way accessible to non-experts, and to highlight outstanding challenges. Along the way we reference where the various derivation of these results, which we mostly omit, can be found in full detail.
@@ 1 Introduction
One particularly interesting limiting case for this setup is the “tensionless” limit, where $T$ is small and the dual theory should become simpler. In fact there is more than one such setup. One is $k=0$ and $h\ll 1$, which can be studied by integrability (by means of the “mirror Thermodynamic Bethe Ansatz” Brollo:2023pkl ), and whose dual should be related to weakly-coupled strings in the D1-D5 system OhlssonSax:2014jtq . Another notable setup is $k=1$ and $h\ll 1$. Here the dual is (a marginal deformation of) the symmetric-product orbifold CFT of four bosons and fermions Giribet:2018ada ; Gaberdiel:2018rqv ; Eberhardt:2018ouy . Also in this case integrability can be identified, both at $h=0$ Frolov:2023pjw and in principle at the leading order in $h\ll 1$ Gaberdiel:2023lco ; Frolov:2023pjw , as we will review. [^2] Note however that it was argued in Fabri:2025rok that the derivation in Gaberdiel:2023lco does not in fact correctly reproduce the integrability structure.
@@ 2.1 Superstring action and integrability
simply by demanding that $J^{(0)}$ and $J^{(2)}$ have eigenvalue $+1$ and $-1$, respectively, under $\Omega$. For a supercoset, one instead needs a $\mathbb{Z}_{4}$ superalgebra automorphism so that the decomposition reads
@@ 2.2 Gauge fixing and lightcone Hamiltonian
It is easy to compute the explicit form of $\mathcal{H}$ as a function of the fields and of their conjugate momenta (see Lloyd:2014bsa for the result including the Fermion contributions). The expression is rather unwieldy (of the Nambu-Goto form) and we do not report here. It is interesting however to look at the expression order by order in the fields. Such an expansion can be obtained as a large-tension expansion ($T\gg 1$, by rescaling the fields by $T^{-1/2}$) and interpreted as a near-pp-wave expansion Berenstein:2002jq .999 Namely, the pp-wave geometry is the one obtained from the metric and B-field described above by expanding around the lightcone geodesic $t(\tau,\sigma)\sim\tau$ and $\varphi(\tau,\sigma)\sim\tau$. At leading order we get
@@ 3 Symmetries, S matrix, and spectrum
The fact that the classical theory underlying $AdS_{3}\times S^{3}\times T^{4}$ strings is integrable gives us hope to quantise the model exactly. The idea is to treat the lightcone gauge-fixed model as one would treat an integrable model like the Sinh-Gordon model. In that case, a powerful approach Zamolodchikov:1978xm is to assume that integrability carries over in the quantum theory and derive consequences of this assumptions. The most dramatic consequences appear for the scattering matrix: the scattering is elastic, without any macroscopic particle production, and any $N$-to-$N$ scattering process can be broken down into a sequence of $N(N-1)/2$ 2-to-2 scattering processes.121212 Note that when we talk about scattering we refer to scattering on the two dimensional worldsheet, where the coordinate $\sigma$ is decompactified, $r\to\infty$. If that is the case, the 2-to-2 scattering matrix is all is needed to completely construct the scattering, and therefore solve the model in the infinite volume limit ($r=\infty$). Along with the usual unitarity requirements, the 2-to-2 S matrix $\mathbf{S}(p_{1},p_{2})$ has to satisfy the celebrated Yang-Baxter equation,
@@ 3.1 Symmetries
In the lightcone gauge-fixed theory we will have fewer symmetries. In terms of the Cartans of $\mathfrak{psu}(1,1|2)_{\mbox{\tiny L}}\oplus\mathfrak{psu}(1,1|2)_{\mbox{\tiny R}}$ we have that the light-cone Hamiltonian141414 To suitably confuse the reader, we indicate the generators of the symmetries in the lightcone gauge with bold letters, and the other generators with regular letters.
@@
Notice that the central extension must vanish on physical states, which obey the level-matching condition $p=0\ \text{mod}2\pi$. This is clearly the case for single-particle states, and on multi-particle states we should get Arutyunov:2006ak ; Borsato:2013qpa
@@
$|Y(p)\rangle_{r}$ $+1$ $+1$ $\tfrac{r}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-2}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ $+1$ $+1$ $\tfrac{r^{\prime}+1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ $+1$ $+1$ $\tfrac{r^{\prime\prime}+2}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|Y(p)\rangle_{r}$ $-1$ $+1$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ $-1$ $+1$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}+1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ $-1$ $+1$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-2}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ $+1$ $-1$ $\tfrac{r-2}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ $+1$ $-1$ $\tfrac{r^{\prime}-1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ $+1$ $-1$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+2}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ $-1$ $-1$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+2}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ $-1$ $-1$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}-1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ $-1$ $-1$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ $+1$ 00 $\tfrac{r-1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-1}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ $+1$ 00 $\tfrac{r^{\prime}}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ $+1$ 00 $\tfrac{r^{\prime\prime}+1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+1}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ $-1$ 00 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+1}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ $-1$ 00 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ $-1$ 00 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-1}{2}$ $\mathbf{1}$ $\mathbf{2}$
@@
$|Y(p)\rangle_{r}$ $+1$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-2}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ $+1$ $\tfrac{r^{\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ $+1$ $\tfrac{r^{\prime\prime}+2}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ $-1$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+2}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ $-1$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}-1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ $-1$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ $+1$ 00 $\tfrac{r-1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-1}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ $+1$ 00 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ $+1$ 00 $\tfrac{r^{\prime\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+1}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ $-1$ 00 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+1}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ $-1$ 00 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ $-1$ 00 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-1}{2}$ $\mathbf{1}$ $\mathbf{2}$
@@
[…]

### v4

@@
11institutetext: Fiona K. Seibold 22institutetext: Blackett Laboratory, Imperial College London, SW7 2AZ London, United Kingdom, and DESY, Notkestrasse 85, 22607 Hamburg, Germany. 22email: fiona.seibold@desy.de (MATRIX Simons Young Scholar) 33institutetext: Alessandro Sfondrini 44institutetext: Dipartimento di Fisica e Astronomia, Università degli Studi di Padova, via Marzolo 8, 35131 Padova, Italy, and Istituto Nazionale di Fisica Nucleare, Sezione di Padova, via Marzolo 8, 35131 Padova, Italy. 44email: alessandro.sfondrini@unipd.it (MATRIX Simons Fellow)
AdS3 Integrability, Tensionless Limits, and Deformations: A Review
Fiona K. Seibold and Alessandro Sfondrini
@@ 1 Introduction
One particularly interesting limiting case for this setup is the “tensionless” limit, where $T$ is small and the dual theory should become simpler. In fact there is more than one such setup. One is $k=0$ and $h\ll 1$, which can be studied by integrability (by means of the “mirror Thermodynamic Bethe Ansatz” Brollo:2023pkl ), and whose dual should be related to weakly-coupled strings in the D1-D5 system OhlssonSax:2014jtq . Another notable setup is $k=1$ and $h\ll 1$. Here the dual is (a marginal deformation of) the symmetric-product orbifold CFT of four bosons and fermions Giribet:2018ada ; Gaberdiel:2018rqv ; Eberhardt:2018ouy . Also in this case integrability can be identified, both at $h=0$ Frolov:2023pjw and in principle at the leading order in $h\ll 1$ Gaberdiel:2023lco ; Frolov:2023pjw , as we will review. [^2] Note that it was argued in Fabri:2025rok that the derivation in Gaberdiel:2023lco presents technical issues that put into question the derivation of the integrability structure.
@@ 2.1 Superstring action and integrability
simply by demanding that $J^{(0)}$ and $J^{(2)}$ have eigenvalue +1+1 and 1-1, respectively, under $\Omega$. For a supercoset, one instead needs a $\mathbb{Z}_{4}$ superalgebra automorphism so that the decomposition reads
@@ 2.2 Gauge fixing and lightcone Hamiltonian
It is easy to compute the explicit form of $\mathcal{H}$ as a function of the fields and of their conjugate momenta (see Lloyd:2014bsa for the result including the Fermion contributions). The expression is rather unwieldy (of the Nambu-Goto form) and we do not report here. It is interesting however to look at the expression order by order in the fields. Such an expansion can be obtained as a large-tension expansion ($T\gg 1$, by rescaling the fields by $T^{-1/2}$) and interpreted as a near-pp-wave expansion Berenstein:2002jq . [^9] Namely, the pp-wave geometry is the one obtained from the metric and B-field described above by expanding around the lightcone geodesic $t(\tau,\sigma)\sim\tau$ and $\varphi(\tau,\sigma)\sim\tau$. At leading order we get
@@ 3 Symmetries, S matrix, and spectrum
The fact that the classical theory underlying $AdS_{3}\times S^{3}\times T^{4}$ strings is integrable gives us hope to quantise the model exactly. The idea is to treat the lightcone gauge-fixed model as one would treat an integrable model like the Sinh-Gordon model. In that case, a powerful approach Zamolodchikov:1978xm is to assume that integrability carries over in the quantum theory and derive consequences of this assumptions. The most dramatic consequences appear for the scattering matrix: the scattering is elastic, without any macroscopic particle production, and any NN-to-NN scattering process can be broken down into a sequence of N(N1)/2N(N-1)/2 2-to-2 scattering processes. [^12] Note that when we talk about scattering we refer to scattering on the two dimensional worldsheet, where the coordinate $\sigma$ is decompactified, $r\to\infty$. If that is the case, the 2-to-2 scattering matrix is all is needed to completely construct the scattering, and therefore solve the model in the infinite volume limit ($r=\infty$). Along with the usual unitarity requirements, the 2-to-2 S matrix $\mathbf{S}(p_{1},p_{2})$ has to satisfy the celebrated Yang-Baxter equation,
@@ 3.1 Symmetries
In the lightcone gauge-fixed theory we will have fewer symmetries. In terms of the Cartans of $\mathfrak{psu}(1,1|2)_{\mbox{\tiny L}}\oplus\mathfrak{psu}(1,1|2)_{\mbox{\tiny R}}$ we have that the light-cone Hamiltonian [^14] To suitably confuse the reader, we indicate the generators of the symmetries in the lightcone gauge with bold letters, and the other generators with regular letters.
@@
[…]
@@
$|Y(p)\rangle_{r}$ +1+1 +1+1 $\tfrac{r}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-2}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ +1+1 +1+1 $\tfrac{r^{\prime}+1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ +1+1 +1+1 $\tfrac{r^{\prime\prime}+2}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|Y(p)\rangle_{r}$ 1-1 +1+1 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ 1-1 +1+1 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}+1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ 1-1 +1+1 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-2}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ +1+1 1-1 $\tfrac{r-2}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ +1+1 1-1 $\tfrac{r^{\prime}-1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ +1+1 1-1 $\tfrac{r^{\prime\prime}}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+2}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ 1-1 1-1 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+2}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ 1-1 1-1 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}-1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ 1-1 1-1 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ +1+1 0 $\tfrac{r-1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-1}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ +1+1 0 $\tfrac{r^{\prime}}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ +1+1 0 $\tfrac{r^{\prime\prime}+1}{2}+\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+1}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ 1-1 0 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+1}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ 1-1 0 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ 1-1 0 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-1}{2}-\tfrac{k}{2\pi}p+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-1}{2}$ $\mathbf{1}$ $\mathbf{2}$
@@
$|Y(p)\rangle_{r}$ - +1+1 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-2}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\eta^{a}(p)\rangle_{r^{\prime}}$ - +1+1 $\tfrac{r^{\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}-1}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|Z(p)\rangle_{r^{\prime\prime}}$ - +1+1 $\tfrac{r^{\prime\prime}+2}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{Z}(p)\rangle_{r}$ - 1-1 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+2}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\bar{\eta}^{a}(p)\rangle_{r^{\prime}}$ - 1-1 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}-1}{2}$ $\mathbf{2}$ $\mathbf{1}$
$|\bar{Y}(p)\rangle_{r^{\prime\prime}}$ - 1-1 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-2}{2}$ $\mathbf{1}$ $\mathbf{1}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ +1+1 0 $\tfrac{r-1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r-1}{2}$ $\tfrac{r}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ +1+1 0 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ +1+1 0 $\tfrac{r^{\prime\prime}+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}+1}{2}$ $\tfrac{r^{\prime\prime}}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|\chi^{\dot{a}}(p)\rangle_{r}$ 1-1 0 $\tfrac{r}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r+1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r}{2}$ $\tfrac{r+1}{2}$ $\mathbf{1}$ $\mathbf{2}$
$|T^{a\dot{a}}(p)\rangle_{r^{\prime}}$ 1-1 0 $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime}}{2}$ $\tfrac{r^{\prime}}{2}$ $\mathbf{2}$ $\mathbf{2}$
$|\tilde{\chi}^{\dot{a}}(p)\rangle_{r^{\prime\prime}}$ 1-1 0 $\tfrac{r^{\prime\prime}}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}-1}{2}+\tfrac{1}{2}\delta H$ $\tfrac{r^{\prime\prime}}{2}$ $\tfrac{r^{\prime\prime}-1}{2}$ $\mathbf{1}$ $\mathbf{2}$
@@
  • A “left” representation has the chiral part of the complex bosons $z,y$ from the transverse modes of the string on $AdS_{3}\times S^{3}$. We denote the excitation created by the related creation operators by $|Y(p)\rangle$, $|Z(p)\rangle$; there are also two fermion modes $|\eta^{a}(p)\rangle$. They all have $\mu=+1$. [^15] Perturbatively, $|{Y}(p)\rangle$ would go to the excitation $a^{\dagger}_{y}(p)|0\rangle$ created by the free oscillator appearing in (24) related to $y$, with $\mu=+1$.